LBC Verification + Collimation model

Observer(s): A. Rakich, O.Kuhn (Tucson), J. Hill, K. Boutsia (LBT)
Telescope Operator: D. Gonzales (LBTO)
Telescope Support:
Instrument Support: J. Morris

Summary

Problems with the dome doors - only the right side is open
PISCES electronics were installed and this task is finished.
Trying to start the telescope, the AzCam was not connected - it took some time to solve this problem.
Dome had to be closed because of clouds, but there have been problems with the latches - it took a long time to close the right door.
The dome shutter system is not robust and we decide not to open for the rest of the night.

Details

02:15 Sky is as good as partly cloudy, and the nearest weather is in Willcox, so we are preparing to open the dome. This is the first time with the new latch pin setup so we have Dan Cox helping with the EPLC on the phone.

until 03:45 Problems with the dome doors - the left side cannot open (see Issue 3588). We only have the right side door open.
  • Since LBCred is not installed, we will go on doing bent gregorian tests (RFBG).

04:03 - Don McCarthy, Craig Kulesa and John M. went up to top off the LN2 and in parallel install PISCES electronics. - this was an afternoon task that was delayed because of J.Little's accident....
  • in the meantime we run some test OBs with LBCblue to verify if the images show in the archive - They do!

05:54 PISCES work is finished, so we prepare to look for a star. Authorize IRTC on Right, none on Left.. Sky is mostly clear.

06:09 Azimuth seems to hit a limit - J.Hill goes up to check.
  • Call D. Ashby - no answer
  • call Dan
  • a switch was in the wrong position - problem seems to be fixed

06:46 GCS problem
  • trying to restart GCS - it takes too long
  • according to the log AzCam is not connected
  • we rebooted the AGW2-azcamg machine
  • the WFS for AZcam machine is not reachable, and we physically can't find it.
  • According to Norm the two cameras are run by the same computer now.
  • GCS seems to be ready - we don't know if it is actullly working

07:17 Closing for clouds
  • problems closing the dome

08:07 Dome closed
  • End of night

-- JohnHill - 04 Sep 2011
Topic revision: r3 - 07 Sep 2011, DavidThompson
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback